# Group-Control Motion Planning Framework for Microrobot Swarms in a Global Field

### v1

@@ 5.2 Instantiations of motion planning problem
For a given set of $n$ robots, we use $m$ group to control it. Depending on the chosen group allocation, the set of all primitives up to order $r_{k}$ is defined as $\mathcal{P}^{r_{k}}=\{f_{j},S^{r_{i}}(\mathcal{F})\}$, where $j=1,...,m$ and $i\leq r_{k}$. This set includes the original $m$ groups and any additional subgroups up to $r_{k}$-th order Lie brackets. One instantiation of the motion planning problem can devise its own primitives set from complete set $\mathcal{P}^{r_{k}}$.
Next, several interesting motion planning instances are discussed to time complexity and optimality trade-offs.
5.2.1 Extreme case 1: Pure planning problem
In this scenario, there are no devised primitives, and the control input is the switch between groups and linear input $u$. Fig.2(a)(d) shows a random run path for allocating 5 robots in 4 groups. The environment setup is a $20\times 20$ square environment with fixed start positions $[0,2;0,4;0,6;0,8;0,9]^{T}$ and goal positions $[4,4;7,7;10,10;12,12;14,14]^{T}$. Circular obstacles are designed in two scenarios with radii 2 and 1, respectively. The path is solved by MATLAB optimization (Fmincon) for a given path length $k=60$, with the cost function of minimizing total control effort $\sum_{i=1}^{k}u$. The random walk collision avoidance algorithm is based on our previous work [17]. This simulation effectively demonstrates that the rotation-only group guides all robots to turn away from obstacles. The simulation is within a minute, but as the number of robots increases, the optimization, as well as random walk obstacle avoidance is no longer time-efficient.
Figure 2: (a)(d)Pure control using Fmincon and random walk obstacle avoidance. (b) Intermediate case 3: RRT using primitives at most order 3 (c), (e) and (f) Intermediate case 4: RRT 3-steps decouple subgroups
5.2.2 Extreme case 2: Pure control problem
The planning problem involves guiding robots one by one using Lie bracket motions. In this case, the primitives are $r_{s}$ order Lie brackets that move only one robot at a time. In this case, the planning is trivial, and what is important is how primitives can be implemented. Two ways of primitives implementation:
  • Repeated Fmincon.
  • Hand-designed primitives.
In the top three figures in Fig.3, Fmincon approximates $[h_{3},g_{1}]$ for a short distance $0.2$ with path length $k=35$. Then, we rotate robots 4 and 5 to their initial orientation. Next, the Fmincon path is repeated, allowing robot 4,5 to continue to move in the goal direction. The repeated path with each length $k=35$ is used to get a path around robots’ neighborhood.
We define the hand-designed primitives as a 3-step control $P_{1}=(f_{i},rot(\pi),f_{i})$, where $f_{i}$ is the vector field of group $i$ in Eq.(5) and $rot(\pi)$ can rotate up to $m$ robots. As the example in Fig.3, $f_{1}=[\mathbf{0}_{6},\cos_{4},\sin_{4},\cos_{5},\sin_{5},\cos_{6},\sin_{6},\mathbf{0}_{6}]$ and $rot(\pi)$ rotate the robot 6 for $\pi$. $P_{1}$ is the primitive for moving robots 4 and 5. In general, if more than $m$ robots need to be rotated, we can use a nested design $P_{j}=(P{j-1},rot(\pi),P{j-1})$.
Figure 3: $[h_{3},g_{1}]$. Top: repeated Fmincon; Bottom: hand-designed primitives
All robots start at the origin and orient to the positive x-axis. Compared with the repeated Fmincon method, nested primitives provide a much simpler path. Also, solving the high dimension of Fmincon faces difficulties. Thus, we state that our designed nested primitives is a better solution.
5.2.3 Intermediate case 3: Controlling rotations
When the order of primitive $r_{k}\geq c_{1}$, the number of robots in a subgroup is no more than $m$, we can control robot angles in that subgroup. Fig. 4(a) gives an example RRT planning using $\mathcal{P}^{c_{1}}$ primitives for 6 robots with 4 groups. The local planner in RRT is to rotate robots in a primitive and then move forward these robots. Initially, robots are deployed in the range $y=[0,10]$ along the $x=0$, and the final states are along $x=10$ with the y-axis positions randomly shuffled. The RRT stops when a goal neighborhood of radius 0.2 is reached. The experiment was repeated 5 times with an average planning time of 72.54 seconds.
5.2.4 Intermediate case 4: Decoupled subgroups
@@
Further increasing the primitive order to $r_{k}=c_{0}$, several primitives can contain a distinct set of robots, and several subgroups together include all robots. Thus, we will plan each distinct subgroup sequentially. This effectively reduces the dimensionality of planning problems. Again, 6 robots with 4 groups are shown in Fig.2(c)(e)(f), where 6 robots were divided into 3 subgroups using Lie brackets $[h2,g1]$, $[h1,g3]$ and $[h3,g2]$. Compared with Fig.2(a). the average planning time shrinks down to 7.44 seconds under the same RRT parameters. Therefore, Decoupled subgroups should be used to deal with large size of swarms.
Next, we demonstrate three levels of motion planning for 4 robots and compare their runtime in Fig.4. Observing the path and runtime data in Tab.3, the running time decreases as we introduce more complex primitives. In particular, the pure planning RRT stops when arriving at the goal neighborhood radius of 2.5, which is far away from the goal, but still takes hundreds of times excessive time compared to the case using a primitive. It shows that RRT does not work unless we introduce subgroups. And there is a balance between robot parallelization and planning time.
(a) Pure planning
(b) Intermediate case: $\mathcal{P}^{2}$
(c) Pure control
Figure 4:
Table 3: Ave Runtime and Tree Nodes for Cases
Case Ave Runtime Ave Tree Nodes Neighborhood radius
Fig.4(a) 35.33 minutes 10583 nodes 2.5
Fig.4(b) 23.46 seconds 148.20 nodes 0.5
Fig.4(c) 4.10 seconds 20.75 nodes 0.5
@@ 6 Conclusion
In conclusion, this paper introduces a novel group-control framework for microrobot swarms in a global field, demonstrating that Small-Time Local Controllability (STLC) is achievable with the minimum number of groups being $log_{2}(n+2)+1$ for n robots. By balancing the complexity of the number of subgroups with motion planning, we have shown that group-control simplifies the motion planning process and scales effectively for larger swarms. Simulations confirm the framework’s effectiveness, suggesting significant potential for applications in various fields requiring precise microrobot coordination. Future work will explore more complex scenarios, larger swarm sizes, and the integration of machine learning methods.

### v2

@@ 6 Instantiations of the motion planning problem
For a given set of $n$ robots, we use $m$ groups to control them. Depending on the chosen group allocation, the set of all primitives up to the order $k$ is defined as $\mathcal{P}^{k}=\bigcup_{i\leq m}\{f_{i}\}\cup\bigcup_{2\leq i\leq k}S^{i}(\mathcal{F})\}$, where $S^{i}(\mathcal{F})$ are subgroups that correspond to Lie brackets of the order $i$. By choosing different subsets of $\mathcal{P}^{k}$, we will obtain different instantiations of the motion planning problem. We next discuss possible implementations of primitives and then investigate several instances of the motion planning problem.
6.1 Implementation of Primitives
Next, we compare two approaches for implementing the primitives: (a) numerical optimization, and (b) primitives designed by hand. In numerical optimization, we choose a large enough number of control steps and randomly decide which group is active during each control step. We then use numerical optimization to compute the time for which each group is activated (the distance for which robots in the group move) so that the total path length is minimized. This approach is similar to the approach described in [17]. In contrast, hand-designed primitives follow the procedure described in the proof of Proposition 4, where robots are eliminated one by one from the original group by using back-and-forth motion.
Figure 2 shows the results in the case of the motion planning problem $M(6,2)$, where we have 6 robots and 4 groups, and the particular motion primitive that we are interested in is the Lie bracket $[h_{2},g_{1}]$ [^1] All the figures indicate the starting location of each robot with a green circle and the goal location with a red cross.. This primitive moves robots 4 and 5 (Table 2). Therefore, the subgroup has 2 robots ($\{4,5\}$), compared to the original group $G_{1}$ that contains 3 robots (4, 5, and 6). The results of numerical optimization are shown on the left. The number of control steps was chosen to be $k=35$. A group is randomly selected for each control step, and then we use numerical optimization to compute the activation time for each group (length of motion) so that the total path length is minimized and that the robots move for some small distance $\epsilon$ in the direction of the target configuration. The distance to the target configuration was $1$, and $\epsilon$ was chosen to be $0.2$; this means that the computed primitive was repeated 5 times to reach the desired goal configuration. The right panel shows the hand-designed primitive using the control vector field $f_{1}=\left[\mathbf{0}_{6},c_{4},s_{4},c_{5},s_{5},c_{6},s_{6},\mathbf{0}_{3},\frac{1}{r},\frac{1}{r},\frac{1}{r}\right]^{T}$ (where $c_{i}\equiv\cos\theta_{i}$ and $s_{i}\equiv\sin\theta_{i}$). The primitive is defined as $P^{1}_{2}=(f_{1},\operatorname{Rot}_{r_{4},r_{5},r_{6}}(0,0,\pi),f_{1})$. The operation $\operatorname{Rot}_{r_{4},r_{5},r_{6}}(0,0,\pi)$ rotates robot 6 by $\pi$ while keeping robots 4 and 5 in place. This rotation involves the rotational vector fields $h_{i}$ as described in Section 4.2. All robots start from the origin and are oriented along the positive $x$-axis. It can be seen that the hand-designed nested primitives in the right panel yield a much simpler path. Moreover, solving optimization problems in high dimensions is challenging.
Figure 2: Motion planning problem $M(6,2)$ with the primitive corresponding to $[h_{2},g_{1}]$ (robots 4 and 5 move, robot 6 moves back-and-forth). Left: numerical optimization. Right: hand-designed primitive.
6.2 Motion planning Cases
Next, we investigate time complexity and efficiency trade-offs for our system’s motion planning problem. Our baseline motion planning algorithm is Rapidly-Exploring Random Trees (RRT) [14]. RRT is a widely used sampling-based algorithm that constructs a tree of feasible trajectories by incrementally extending it from an initial state toward randomly sampled points in the configuration space. The tree grows by iteratively sampling the points (nodes) and connecting them to the nearest node in the existing tree using a local planner. In our formulation, the local planner uses subgroups. In some cases, it further simplifies the problem by rotating the robots in the subgroup so they are directed toward the sampled points before using the motion primitive corresponding to the subgroup.
a    b    c
Figure 3: EC1: (a) Numerical optimization (exact solution); (b) RRT with rotation; (c) Original RRT.
Algorithm Runtime RRT Nodes Path Length Execution
Numerical Opt. 0.80s NA 181.04 76.33s
RRT with rot. 8.88s 147.20 323.57 172.36s
Original RRT2 393.53s 7736.80 332.10 110.70s
Table 3: EC1 Algorithm Performance (average of 20 instances). Algorithm 1 is exact, Algorithms 2 and 3 terminate within a neighborhood of radius 2. 2No solution was found for Original RRT in 15 out of 20 cases within the allotted time.
6.2.1 Extreme Case 1 (EC1): Pure planning
In this scenario, there are no additional subgroups; only the original groups are used. Thus, each robot belongs to more than one group, which means the robot motions are coupled. Figure 3 compares numerical optimization (see Section 6.1) and RRT planning in the case of 4 groups and 6 robots. Numerical optimization does not consider collisions but finds the path that reaches the target exactly. Both RRT with rotation (Fig. 3b) and Original RRT (Fig. 3c) use the original groups as local planners and find a path that reaches the target within a certain neighborhood radius. The difference is that RRT with rotation aligns the robots with the intermediate goal point before moving them, while Original RRT does not, so it searches in higher-dimensional space. Initially, the robots are deployed in the range $y=[0,10]$ along $x=0$, and the final goal positions are along $x=15$, with the $y$ coordinates permuted randomly. Specifically, robots have start positions $[(0,2.0),(0,5.2),(0,8.4),(0,11.6),(0,14.8),(0,18.0)]^{T}$ and goal positions $[(12,11.6),(12,18.0),(12,5.2),(12,2.0),(12,8.4),(12,14.8)]^{T}$. The density of the robot icons on the path shows the efficiency of the motion. Numerical optimization provides an efficient path and icons are sparse. RRT with rotation has clusters of icons on the path; each cluster illustrates how robots move when we only control the orientations of some robots. Pure RRT has a high density of icons on the path, which means it is less efficient. Table 3 compares the performance of these three algorithms in terms of computation, path length, and execution time. The results are reported for an average of 20 trials. Numerical optimization is the fastest and produces the shortest path. Unfortunately, it becomes computationally prohibitive for collision avoidance and in complex environments. As for the other two algorithms, RRT with rotation has a much faster runtime compared to RRT, but it incurs a penalty in the execution time.
a (EC2) b (EC1) c (IC3)
d (IC4 4, 5) e (IC4 1, 3) f (IC4 2, 6)
Figure 4: (a) EC2: Pure control; (b) EC1: Pure planning using RRT with rotation; (c) IC3: Parallel motion; (d) IC4: Robots 4 and 5; (e) IC4: Robots 1 and 3; and (f) IC4: Robots 2 and 6.
Case RRT Runtime RRT Nodes Path Length Execution
RRT with rotation3 1141.38s 8242.75 384.62 200.57s
Subgroup parallel 216.15s 2218.01 868.66 886.04s
Subgroup sequential 15.38s 102.10 529.47 577.41s
Pure control 19.44s 175.50 871.82 1042.47s
Table 4: Performance of the four motion planning algorithms with obstacles (average of 10 instances). All algorithms terminate within a neighborhood of radius 1.5. 3No solution was found for Original RRT in 8 out of 10 cases within the allotted time.
6.2.2 Extreme Case 2 (EC2): Pure control
In this scenario, each subgroup contains a single robot. The planning problem involves guiding individual robots to the goal. The primitives are of order $r=3$ and they move one robot at a time. Figure 4a shows a sample path for 6 robots in 4 groups. The environment is the square $[0,20]\times[0,20]$ containing two black circular obstacles. The icon density is high on the path, and each repeated icon pointing in the same direction indicates the back-and-forth motion in primitives.
6.2.3 Intermediate Case 3 (IC3): Parallel planning
In this case, we explore the trade-off between the path execution time and RRT planning computation time. Figure 4c presents an example in which we use $3$ disjoint subgroups, each containing $2$ different robots. These subgroups are inspired by the Lie brackets $[h_{2},g_{1}]$, $[h_{3},g_{2}]$ and $[h_{1},g_{3}]$. This experiment was repeated $10$ times, with an average planning time of $216.15$s and an execution time of $886.04$s.
6.2.4 Intermediate Case 4 (IC4): Sequential planning
@@
In contrast to IC3, here we plan the motion of each subgroup sequentially. This effectively reduces the dimensionality of planning problems. Figures 4d-f show the motions where 6 robots were divided into 3 subgroups using Lie brackets $[h_{2},g_{1}]$, $[h_{1},g_{3}]$ and $[h_{3},_{2}]$. Compared with Fig. 4c, the average planning time, path length, and execution time decrease to $19.44$s, $529.47$, and $577.41$s, respectively, under the same RRT parameters.
Table 4 compares the performance of different strategies for solving the motion planning problem in the presence of obstacles. RRT generates the most efficient path, but it takes considerably longer to find it. On the other hand, pure control is very efficient in finding the path, but the path is inefficient. It can be seen that the sequential plan provides an excellent compromise between the computation time and the efficiency of the generated path. For large swarms, sequential planning is computationally much more efficient than parallel planning.
@@ 7 Conclusion
This paper investigates how a novel group-control framework can be used for motion planning for microrobot swarms controlled by a global field. We focus on a swarm of MicroStressBots, electrostatically actuated MEMS stress microrobots. We show that the system is Small-Time Locally Controllable (STLC) in positions even though the control inputs are unilateral. We also demonstrate that the minimum number of groups to achieve STLC for $n$ robots is $log_{2}(n+2)+1$. We study the complexity of the motion planning problem for a MicroStressBot swarm under group-based control. In particular, we compare the trade-off between the complexity of control, the complexity of motion planning, and the efficiency of the generated path. We introduce the notion of motion primitives and subgroups, which allows us to balance these different factors. We show that subgroups significantly simplify the motion planning problem and make the approach applicable to larger swarm sizes. Simulations confirm the effectiveness of the framework, suggesting significant potential for applications requiring microrobot coordination, such as microassembly, tissue engineering, and drug delivery.